LINEAR SCALING DENSITY FUNCTIONAL METHODS FOR EXCITATION ENERGIES AND VIBRA-TIONAL FREQUENCIES OF LARGE MOLECULES

\underline{GUSTAVO E. SCUSERIA}, Chemistry Department, Rice University, Houston, TX 77251-1892.

Thie presentation will address recent advances in electronic structure methods for large molecules. The focus will be on linear scaling density functional theory (DFT) methods. Methodological developments will be discussed and applications to excitation energies and vibrational frequencies of large molecules will be presented.